SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (26th June, 2007)

Attendees: Debbie Bard, Rainer Bartoldus, Al Eisner, Kim Hojeong, Rahmat Rahmat, Mark Tibbetts.

Operations

  • Recieved email from DQM over the weekend about GLT JAS instructions. They've been clarified. Also made a note on L3 JAS page that L3 EMC plot is rotated compared to the EMC system plots.

EMT

  • Broken board still with Lupe.
  • New EMT cellphone arrived - the number stays the same.

DCT/DCZ

  • Problems over the weekend with the DCH - see hypernews discussion. They will replace the electronics box tomorrow.
    • Should check TRG plots to see if this has caused any problems in the DCT.
    • They have not masked out any layers yet but may have to today if the electronics fail completely. In this case, we'll not have enough layers for a B track. We can alter the BLT mask to set the relevent bit to 1 which will regain efficiency at expense of backgrounds.
    • Should perform a DCH-DCT calibration after the work is done.

L3

  • MC validation of new L3 tracking still up for grabs. It really should be implemented by the end of this Run.
  • Hojeong has been having trouble with L1TNtuple - see hypernews posting.

Data Quality and Monitoring

  • Run 6 TRG stripcahrts are ~90% done.
  • Rahmat is in the process of obtaining an account at Padova.
  • The DQG has noticed the pi0 mass goes down as the backgrounds increase. This is curious.

Trigger Configuration

  • The study still needs validaation on SP9 MC.
  • The propsed 1E trigger line requires a re-think of L3 flags. The BGFilters are nearly ready, so L1 and L3 will soon be the issues preventing implementation.

This page is created by Debbie Bard (djbard@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.